Bus Econ 13.5.69 Question Help A company manufactures mountain bikes. The research department produced the marginal cost function C(x)-500-งิ 0sxs900 where C(x) is in dollars and x is the number of bikes produced per month. Compute the increase in cost going from a production level of 600 bikes per month to 900 bikes per month. Set up a definite integral and evaluate it. The increase in cost is s
Bus Econ 13.5.81 := Question Help Given the supply function p=S(x) = 7(-0.02x-1) find the average price (in dollars) over the supply interval [26,33]. The average price is $ (Type an integer or decimal rounded to two decimal places as needed.)
Answer #1

1)

we have

5 C,(z) = 500-2

increase in cost from a production level 600 to 900,

900 600

900 C(x)-/ 500-_d.r

C(z) = 1500r-6 C(z) | 5002

(600)2 6- (900% _ 500(600) 500(900) C(z) --

C(x) 75000

2)

we have

p=7(e^{0.02x}-1)

the average value is,

Pavg b-a

put the value of S(x), a = 26 and b = 33 in above equation,

r33 Pavg33 - 26 J26

702d J25 1 p aug 7

r33 0.02r - 1)dx (e Pavg J26

p_{avg}=\left [ \frac{e^{0.02x}}{0.02}-x \right ]_{26}^{33}

p_{avg}=\left [50e^{0.02x}-x \right ]_{26}^{33}

50e002(33) 33 50e 0.02(26) +26

p_{avg}=5.63823